Find f'(x) of f(x)= sin²x.sinx².sin²x²


1
Expert's answer
2021-03-03T04:21:15-0500


If f1(x) = sin²x.sinx² and f2(x) = sin²x² then

f(x) = f1(x)*f2(x)

f'(x) = f1'(x)*f2(x) + f1(x)*f2'(x)= (sin²x.sinx²)’.sin²x²+ sin²x.sinx².(sin²x²)’

f1'(x) = (sin²x.sinx²)’ = 2sinx cosx sinx2+sin2x 2xcosx2 = 2sinx (cosx sinx2 +x sinx cosx2)    

f2'(x) = (sin²x²)’ = 2 sinx2 cosx2 2x = 4x sinx2 cosx2

f'(x) = 2sinx (cosx sinx2 +x sinx cosx2) sin²x²+ 4x sinx2 cosx2 sin²x sinx² =

= sin²x² [2sinx (cosx sinx2 +x sinx cosx2)+ 4x cosx2 sin²x] =

= 2sinx sin²x² [cosx sinx2 +x sinx cosx2+ 2x cosx2 sinx] =

= 2sinx sin²x² [cosx sinx2 +3x sinx cosx2]

Answer:

f'(x) = 2sinx sin²x² [cosx sinx2 +3x sinx cosx2]